OpenStudy (anonymous):

I know this is simple, that probably why I am unsure of the answer. d/r+d-d/r-d=

OpenStudy (anonymous):

\[\frac{d}{r+d}-\frac{d}{r-d}????\]

OpenStudy (anonymous):

\[\frac{d}{r+d}-\frac{d}{r-d}= d(\frac{1}{r+d}-\frac{1}{r-d})\] : P

OpenStudy (anonymous):

subtract and simplify

OpenStudy (anonymous):

\[\frac{d(r-d)-d(r+d)}{(r-d)(r+d)}\]

OpenStudy (anonymous):

\[\frac{-2d^2}{(r-d)(r+d)}\]

OpenStudy (anonymous):

\[\frac{d(r-d)-d(r+d)}{(r+d)(r-d)}\] \[\frac{dr-d^2-dr-d^2}{r^2-d^2}=\frac{-2d^2}{r^2-d^2}\]